If we have more than one vectors ,how can we add? Say we have vectors a,b,c and d.

Grade:11

1 Answers

BALAJI ANDALAMALA
8 years ago
Let a and b are two vectors .

a= xi+yj
b=pi+qj

While adding , x components should be added to x-components only.

a+b = (x+p)i+(y+q)j
Follow the same procedure in case of four vectors also.